OpenStudy (anonymous):

Choose the replacement for w that makes the equation a true statement. -3w – 24 = -39 A.5 B.-5 C.21 D.-21

pooja195 (pooja195):

Plug in each of your options to see if the statement is true

pooja195 (pooja195):

A. -3(5)– 24 = -39 B. -3(-5) – 24 = -39 C. -3(21) – 24 = -39 D. -3(-21)– 24 = -39 From here just solve all of these and see which one is a true statement

OpenStudy (purplemexican):

-3(5)-24=-39